Make New Entry,
Make Followup Entry
User name Ramesh
Log entry time 14:55:31 on May 8,2011
Entry number 358621
keyword=Right hrs to 24.5 deg, ramping up
We (Howard Smith and myself) moved the right spectrometer from 28 deg to 24.5 degrees. We are also ramping up this spectrometer from zero (while it was made ready for polarity change) to 1.19635 GeV/c. Also cycled Q2 to 1400 Amp and Q3 to 1100 Amp. Dipole is ramping up as well.